OpenStudy (anonymous):

The answer is y = 4, right? What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 -3x + 2y = 1 2 4 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (amriju):

yup..

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

Can you help me with this one? I can't remember how to solve it. :) What is the value of the x variable in the solution to the following system of equations? 3x + y = 6 6x + 2y = 8 zero 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (amriju):

ok..so what u do basically is multiply 2 on both sides of eqn 1...(the first eqn)...then subtract the two eqns...u will get 0x=4..impling no solution the system...since x would be 4/0 thats infinity...

OpenStudy (amriju):

basically if u see u'll find that 6x+2y=12 ( dividing by 2 on both sides will give u eqn 1)..and 6x+2y=8..clearly thats not possible..

OpenStudy (anonymous):

Oh, ok... So would x = 4, and y = -6?

OpenStudy (amriju):

no..i think it wud be 7/3

OpenStudy (anonymous):

@Niall Yes. Add the two equations LHS to LHS and RHS to RHS. The result is 3y = 12.

OpenStudy (amriju):

if u mean the first ques u posted..

OpenStudy (anonymous):

I'm gonna post some screen shots of my problems. Can someone check them for me and make sure that I have all of the correct answers?

OpenStudy (amriju):

just ques 4 is wrong i think...the others r correct..u posted ques 4..